NI 20Hz bump investigation

Go down

NI 20Hz bump investigation

Post  Alessio on Wed Jul 22, 2009 5:24 pm

As stated in some entries in the Virgo Logbook, it is clear that the NI bench has a non negligible coherence with dark fringe around 20-30Hz. A bump is clearly visible in the NI CO2 laser noise spectrum. With the idea that the bump could be due to an acoustic coupling with the fans of the RF driver placed beneath the NI bench, we moved the driver to the TCS room.
When we turned on again the laser, power on the PD was above the saturation level of the pre-amplifier for no apparent reason. Turning the waveplate had no effect. We changed the chiller setpoint by 0.3 degrees and TCS power came back to normal values (see fig. 1).
We could than take a noise spectrum to see the effect of moving the RF driver outside the acoustic enclosure (see fig. 2). As it is evident, there is no change at all, the bump at 20-30Hz is still there.
Since we changed the chiller setpoint, thus the laser body temperature, we checked that the shape of the ring had not changed. Figure 3 shows the two rings as measured by the IR cameras. Comparing with the old images, it seems there is no difference at all.





avatar
Alessio

Posts : 127
Join date : 2008-11-07
Age : 46
Location : Rome

View user profile

Back to top Go down

Back to top

- Similar topics

 
Permissions in this forum:
You cannot reply to topics in this forum